OpenStudy (anonymous):

Help needed please for exams Find cot60 degrees-sin45degrees a)2-sqrt3/2 b.)2 square root3-3 square root 2/6 c)2-squareroot2/2 d)2 square root 2-3 square root 3/637 I got 1/ square root 3 -1/ square root 2 ...what is done next?

OpenStudy (amistre64):

it would be better if you do not post this as an "exam" question :/

OpenStudy (amistre64):

your values are fine, you might want to move up the sqrts tho; then its just subtracting fractions

OpenStudy (anonymous):

its a practice past exam paper and i dont understand it.its not being graded

OpenStudy (amistre64):

\[\frac{\sqrt3}{3}-\frac{\sqrt2}{2}\]

OpenStudy (amistre64):

you prolly have forgotten, or dont remember how to move a radical consider this: \[\frac{1}{\sqrt a}\] \[1*\frac{1}{\sqrt a}\] \[\frac{\sqrt a}{\sqrt a}*\frac{1}{\sqrt a}\] \[\frac{\sqrt a}{(\sqrt a)^2}\] \[\frac{\sqrt a}{a}\]